# On Variational Perspectives To The Graph Isomorphism Problem

### v2

@@ Abstract
We consider a quadratic unconstrained binary optimization (QUBO) formulation of the graph isomorphism problem from a variational quantum algorithmic perspective. By treating it using the Quantum Approximate Optimization Algorithm (QAOA) and the Variational Quantum Eigensolver (VQE), this study presents results for each and highlights variations therein for graphs of four and five nodes. For isomorphic graphs having an equal number of nodes and edges, we demonstrate clustering in the energy landscape for the QAOA. This trend found in the QAOA study was further reinforced by studying the ground state energy reduction using VQEs. Furthermore, we examine the trend under which isomorphic pairs of graphs vary in the ground state energies, with varying edges and nodes.
@@ I Introduction
What underpins quantum annealing is the adiabatic theorem in physics, which involves slowly evolving the ground state of a physical system from the problem Hamiltonian of the system, and is what enables finding the ground state energies relevant to QUBO problem formulations Santoro and Tosatti (2006); Lewis and Glover (2017); Domino et al. (2021). Most prior quantum information efforts to tackle the graph isomorphism problem have focused on quantum annealing approaches. While annealers are effective for specific instances, they face limitations in scalability, connectivity, and flexibility. In contrast, VQAs provide greater adaptability to different problem formulations on account of relying on a gate-based circuit model. They are furthermore empowered by the hybrid quantum-classical optimization loop, making them more robust for implementations on noisy intermediate-scale quantum (NISQ) devices. By considering VQAs for tackling graph isomorphism, this work explores whether the variational paradigm offers not only practical performance on small graph instances, but also a potential pathway to scalable quantum heuristics for one of the most enigmatic problems in complexity theory.
Prior QUBO formulations for the Graph Isomorphism Problem have been provided by LucasLucas (2014), Kenneth Zick, Shehab, and French (2015), Calude Calude, Dinneen, and Hua (2017) etc. Kenneth and Calude performed their QUBO formulation experiments on D-Wave’s Quantum Annealer machines. It has been conjectured that the lowest number of logical qubits required for the general Graph Isomorphism Problem for two graphs of order n is $n^{2}$. Apart from this, leveraging QUBO for a problems such as Boolean satisfiability Farhi et al. (2000), binary clustering Bauckhage et al. (2017), classifier training Pudenz and Lidar (2013), Travelling Salesman Problem Salehi, Glos, and Miszczak (2021) vehicle routing Borowski et al. (2020), railway dispatching problem Domino et al. (2021) has been explored. This work presents a first proof-of-principle approach to solving the Graph Isomorphism Problem using the circuit model of quantum computation through variational means in a NISQ paradigm.
@@ II.1 The Quantum Approximate Optimization Algorithm
$\lvert\mathbf{\alpha},\mathbf{\beta}\rangle=\prod_{k=1}^{L}\textrm{exp}(-i\beta_{k}H_{\text{m}})\textrm{exp}(-i\alpha_{k}H_{\text{o}})\lvert+\rangle^{\otimes n},$ (1)
where $H_{\text{o}}$ is the objective and $H_{\text{m}}=-\sum_{j}X_{j}$ is the mixer Hamiltonian. Under fixed number of layers (i.e. $L$), the algorithm requires $2L$ parameters. The expectation value $\langle\alpha,\beta|H|\alpha,\beta\rangle$, of state $\lvert\alpha,\beta\rangle$ is approximated through measuring the state in the computational basis. The parameters $\alpha$, $\beta$ are updated using classical optimization procedures so that the energy returns an optimal value. QAOA can be harnessed to solve problems such as Max-CutFarhi et al. (2001); Majumdar et al. (2021), Max k-Vertex CoverCook, Eidenbenz, and Bartschi (2020) and even integer factorization Anschuetz et al. (2019).
@@ II.2 The Variational Quantum Eigensolver
$\lvert\psi(\mathbf{\theta})\rangle=U(\mathbf{\theta})\lvert 0\rangle,$
@@ III Formulation of quadratic program
$\begin{split}H=A\sum_{v}\left(1-\sum_{i}x_{v,i}\right)^{2}+A\sum_{i}\left(1-\sum_{v}x_{v,i}\right)^{2}+B\sum_{\begin{subarray}{c}ij\notin E_{1}\\ i\neq j\end{subarray}}\sum_{uv\in E_{2}}x_{u,i}x_{v,j}+B\sum_{ij\in E_{1}}\sum_{\begin{subarray}{c}uv\notin E_{2}\\ u\neq v\end{subarray}}x_{u,i}x_{v,j}.\end{split}$ (3)
@@ IV.1 Numerical Setup for QAOA
The Quantum Approximate Optimization Algorithm (QAOA) was implemented in Qiskit using the QAOA class. The quadratic program, formulated following the approach of Calude et al. Calude, Dinneen, and Hua (2017), was first expressed in IBM’s Docplex framework, which enabled its direct translation into the corresponding problem Hamiltonian. This Hamiltonian was then supplied to the QAOA class through the minimum eigensolver interface, which evaluates the minimum eigenvalue associated with the encoded cost function. We restrict to p = 1 to remain in the shallow-circuit regime typical of NISQ demonstrations, balancing fidelity with feasibility.
@@ IV.3 Limitations
Our experiments were restricted to graphs of order four and five due to both hardware and algorithmic constraints. In the QUBO formulation, the number of logical qubits scales as $\mathcal{O}(n^{2})$ with the number of vertices $n$, making even moderately larger instances infeasible on current NISQ devices with limited qubit counts, connectivity, and coherence times. Classical simulation is likewise hindered by the exponential growth of the Hilbert space. Further limitations arose from algorithmic overhead. The VQE, even with efficient optimizers such as SPSA, is challenging to scale to larger graphs, and remains resource-intensive. For these reasons, we confined our study to small graphs in order to establish proof-of-principle results. This choice enabled us to demonstrate energy clustering for isomorphic graph families under controlled conditions, while leaving the extension to larger and more complex graphs to future work. Progress in compact QUBO encodings, problem-tailored ansätze, and error mitigation techniques will be necessary to overcome these current limitations.
V Results and Discussion
V.1 QAOA Results
The energies calculated are of the minimum eigenvalues of the converted problem Hamiltonian from the initial quadratic Docplex model using the MinimumEigenOptimizer class that takes the quadratic model and converts them into a tensor product of Pauli matrices. The corresponding minimum energy eigenvalues of the Hamiltonian for which the quadratic program yields a minimum eigenvalue of zero are given in Fig.˜1(a). Similarly, for graphs with five nodes, the energies are presented in Fig.˜1(b). Extending on the results by Szegedy Szegedy (2019), in the light of the Ising formulation Bonati (2014); Istrail (2000); Barahona (1982) of the Graph Isomorphism Problem, solved using the QAOA, we observe that the energies for graphs having the same node count have roughly the same energies, as is evident from the graphs. Moreover, the number of edges have roughly the same energies, with the energies values being distributed in a space having a relatively small value of standard deviations.
Szegedy’s result Szegedy (2019) states that for a MAX-CUT formulation of QAOA, two isomorphic graphs will yield the same energy values given the same parameters for the QAOA quantum circuit. Our results support this by showing that families of isomorphic graphs have energies within a very short deviation range. We keep formalising this argument more carefully for future work. However, in an experiment conducted as part of this study that takes two identical complete graphs and attempts to methodically remove one edge at random without replacement, varying the number of edges and the number of nodes in the graph simultaneously yields the results given in Fig.˜2. Interestingly, as edges and nodes are reduced, the energy decreases up to edge number 6, but then rises again for sparser graphs. This non-monotonic trend may reflect finite-size effects, ansatz expressibility limits, or optimizer artifacts, and warrants further investigation and study.
@@ V.1 QAOA Results
Figure 2: The illustration of variation in QAOA energies with simultaneous edge and node reduction. (a) Shows how edges were methodically removed from a five node complete graph (the graphs are reduced from (j) $\rightarrow$ (a)). A copy of each graph was used when testing, thereby yielding a pair of identical graphs, on which the graph isomorphism formulation was carried out. (b) Shows the QAOA energies of the Hamiltonian for each of these graphs for which the isomorphic pairs has the quadratic program yield a function value of zero
V.2 VQE Results
The findings obtained from QAOA were further corroborated by the results of the VQE. In particular, VQE yielded an optimal function value of zero for isomorphic graph pairs and a non-zero value for non-isomorphic pairs, in agreement with the theoretical expectations of the QUBO formulation. This outcome demonstrates that the variational principle underlying VQE is able to faithfully capture the structural equivalence encoded in the Hamiltonian, thereby validating the observations made with QAOA. A notable feature of the VQE results is that its ground-state energies were of much larger magnitude than those from QAOA. This stems from VQE’s deeper ansatz and broader variational search, which approximate the ground state with higher fidelity.
Although the absolute energy scales differ, both methods showed consistent trends: isomorphic graph families clustered tightly in energy, while non-isomorphic pairs deviated. This clustering thus emerges as a variational signature of graph isomorphism and supports the use of energy landscapes as a heuristic for distinguishing graph families. It is important to emphasize, however, that this improved fidelity comes at a substantial computational cost. Unlike QAOA, which requires only a shallow circuit and a limited parameter set, VQE relies on repeated circuit preparation and measurement at each optimization step. In our implementation, this involved up to 600 iterations of the SPSA optimizer on a TwoLocal ansatz, resulting in significantly longer execution times and higher resource usage. This trade-off highlights a key distinction between the two approaches: QAOA offers a lightweight variational heuristic capable of rapidly identifying clustering patterns, whereas VQE provides a more resource-intensive but rigorous validation of those patterns. The convergence of VQE energies across iterations for multiple runs on four-node graph pairs is shown in Fig.˜3.
Taken together, these results suggest a complementary role for the two algorithms. QAOA, with its shallow-depth circuits, can serve as a practical diagnostic tool for detecting candidate isomorphic families, while VQE, through Ritz’s variational principle, can confirm these findings. The fact that both methods independently reproduce the clustering phenomenon strengthens the central claim of this work: that variational algorithms provide a viable pathway toward quantum heuristics for the Graph Isomorphism problem, even in the small-instance regime accessible on near-term devices. Furthermore, the consistency between QAOA and VQE supports the broader hypothesis that energy clustering is not an artifact of a particular ansatz or optimization method, but rather an intrinsic property of the Hamiltonian encoding of isomorphism.

### v3

# Energy Landscape Structure of Small Graph Isomorphism Under Variational Optimization

@@ Abstract
We investigate a quadratic unconstrained binary optimization (QUBO) formulation of the graph isomorphism problem using the Quantum Approximate Optimization Algorithm (QAOA) and the Variational Quantum Eigensolver (VQE). For small graph instances, we observe that isomorphic pairs exhibit consistent clustering in variational energies, indicating that the Hamiltonian successfully encodes structural features. However, we demonstrate that low variational energy alone is an unreliable certifier of isomorphism due to the high probability of converging to infeasible states that violate bijection constraints. To address this, we analyze optimization trajectories rather than final energies; consistently outperform naive energy thresholding, though absolute performance remains limited. Our results characterize the current limits of variational algorithms for graph isomorphism, positioning energy landscape analysis as a diagnostic tool rather than a scalable decision procedure in the NISQ regime.
@@ I Introduction
What underpins quantum annealing is the adiabatic theorem in physics, which involves slowly evolving the ground state of a physical system from the problem Hamiltonian of the system, and is what enables finding the ground state energies relevant to QUBO problem formulations Santoro and Tosatti (2006); Lewis and Glover (2017); Domino et al. (2021). Most prior quantum information efforts to tackle the graph isomorphism problem have focused on quantum annealing approaches. While annealers are effective for specific instances, they face limitations in scalability, connectivity, and flexibility. In contrast, VQAs provide greater adaptability to different problem formulations on account of relying on a gate-based circuit model. They are furthermore empowered by the hybrid quantum-classical optimization loop, making them more robust for implementations on noisy intermediate-scale quantum (NISQ) devices. This work presents an exploratory small-instance, proof-of-principle study within a NISQ framework, complementing earlier quantum annealing approaches and highlighting optimization behaviour and limitations. While quantum annealing has proven effective for small instances of graph isomorphism, it faces several intrinsic limitations. First, annealers rely on fixed hardware connectivity graphs, requiring costly minor embedding for dense or irregular QUBO instances, which can significantly inflate qubit overhead and degrade solution quality. Second, the annealing schedule and mixer dynamics are largely fixed, limiting flexibility in exploiting problem-specific structure or symmetries. Third, annealers primarily provide access to the ground state, offering limited insight into the broader energy landscape that may encode useful structural information.
Gate-based variational algorithms address these limitations by enabling fully programmable Hamiltonians, mixers, and ansätze, allowing greater flexibility in encoding constraints and symmetries. Moreover, variational methods provide access not only to ground-state energies but also to the structure of low-energy excitations and energy clustering phenomena. This flexibility is particularly valuable in the NISQ era, where error mitigation, adaptive circuit design, and hybrid quantum-classical workflows can be tailored to specific problem instances. As circuit-model quantum hardware continues to scale, variational approaches offer a natural pathway toward more expressive and scalable quantum heuristics for graph isomorphism.
By considering VQAs for tackling graph isomorphism, this work explores whether the variational paradigm offers not only practical performance on small graph instances, but also a potential pathway to scalable quantum heuristics for one of the most enigmatic problems in complexity theory.
I.1 Related work
Quantum approaches to the Graph Isomorphism (GI) problem have historically followed two largely distinct directions. The first line of work attempts to cast GI as an instance of the non-Abelian Hidden Subgroup Problem (HSP), solvable in principle via quantum Fourier sampling. Early optimism surrounding this approach was tempered by a series of negative results demonstrating that standard coset-state and Fourier-sampling techniques fail to efficiently distinguish automorphism groups, even for restricted families of graphs Hallgren et al. (2010); Grigni et al. (2001); Shehab and Jr. (2017). As a consequence, the HSP-based approach is now widely regarded as unlikely to yield a general and efficient quantum algorithm for graph isomorphism.
A second, more practically successful direction encodes GI as a quadratic unconstrained binary optimization (QUBO) or Ising Hamiltonian, solvable via quantum annealing. Seminal works by Lucas Lucas (2014), Zick et al. Zick, Shehab, and French (2015), and Calude et al. Calude, Dinneen, and Hua (2017) demonstrated that the GI problem can be mapped to an energy minimization task in which valid isomorphisms correspond to ground states of a suitably constructed Hamiltonian. These formulations have become the dominant quantum heuristic for GI and align with the broader success of QUBO-based models for combinatorial optimization problems such as Boolean satisfiability and the Traveling Salesman Problem Farhi et al. (2000); Salehi, Glos, and Miszczak (2021).
Despite their effectiveness, annealing-based approaches face intrinsic hardware-level limitations. The dense logical connectivity required by GI QUBOs must be embedded into sparse physical hardware graphs, often necessitating complex minor-embedding procedures. This embedding overhead substantially inflates the number of required physical qubits relative to the logical problem size, thereby limiting scalability on current annealing platforms. Furthermore, standard annealers employ fixed driver Hamiltonians, restricting the flexibility with which the energy landscape can be explored.
In contrast, comparatively little work has investigated graph isomorphism using gate-based variational quantum algorithms (VQAs), such as the Quantum Approximate Optimization Algorithm (QAOA) or the Variational Quantum Eigensolver (VQE). To the best of our knowledge, there are no prior demonstrations applying such circuit-model variational methods directly to the GI problem. This absence is largely explained by resource considerations: a natural QUBO encoding of GI for graphs of order $n$ requires $O(n^{2})$ logical qubits, a scale that has only recently become accessible for small instances on gate-based quantum hardware and high-fidelity simulators. Historically, annealing architectures provided the only platforms with sufficient qubit counts to explore such encodings, even at modest problem sizes.
The motivation for revisiting GI within a gate-based variational framework is therefore complementary rather than competitive with quantum annealing. While annealers excel at large-scale ground-state search within fixed Hamiltonian families, variational circuit-model approaches offer distinct advantages in terms of flexibility and analytical access. Logical connectivity can be handled via compilation techniques, such as SWAP networks, rather than physical embedding, potentially mitigating qubit overhead at the expense of circuit depth. Moreover, variational algorithms allow for programmable ansätze, non-commuting mixer Hamiltonians, and access to intermediate optimization dynamics, features that are largely inaccessible in standard annealing paradigms. We emphasize that this work does not claim a computational advantage over classical algorithms for graph isomorphism, but instead focuses on characterizing the behavior and limitations of variational quantum algorithms applied to isomorphism-encoding Hamiltonians in the small-graph regime.
Recent theoretical results further motivate this perspective. Szegedy Szegedy (2019) showed that QAOA expectation values at fixed depth define graph invariants, suggesting that variational energy landscapes may encode nontrivial structural information beyond strict ground-state optimization. This observation motivates an exploratory study of how graph structure manifests in variational energy landscapes, even when exact isomorphism certification is infeasible.
Accordingly, the present work should be viewed as a proof-of-principle investigation of graph isomorphism within the gate-based variational paradigm, restricted to small graphs in the NISQ-accessible regime. Rather than proposing a scalable solution to GI, we aim to characterize how isomorphism-encoding Hamiltonians behave under variational optimization, to identify both the capabilities and limitations of energy-based heuristics, and to clarify the role of variational dynamics in extracting structural information from quantum encodings of GI.
@@ II Background
In light of these negative results regarding the standard algebraic frameworks, recent attention has shifted towards heuristic optimization methods better suited for Near-Term Intermediate Scale Quantum (NISQ) devices. The following sections detail the theoretical foundations of the Quantum Approximate Optimization Algorithm (QAOA) and the Variational Quantum Eigensolver (VQE), which serve as the primary variational architectures for our proposed approach.
@@ II.1 The Quantum Approximate Optimization Algorithm
[…]
@@ II.2 The Variational Quantum Eigensolver
[…]
@@ III Formulation of quadratic program
[…]
@@ IV.1 Numerical Setup for QAOA
The Quantum Approximate Optimization Algorithm (QAOA) was implemented in Qiskit using the QAOA class. The quadratic program, formulated following the approach of Calude et al. Calude, Dinneen, and Hua (2017), was first expressed in IBM’s Docplex framework, which enabled its direct translation into the corresponding problem Hamiltonian. This Hamiltonian was then supplied to the QAOA class through the minimum eigensolver interface, which evaluates the minimum eigenvalue associated with the encoded cost function. We restrict to $p=1$ to remain in the shallow-circuit regime typical of NISQ demonstrations, balancing fidelity with feasibility. Here, $p$ denotes the number of QAOA layers, with each layer consisting of one application of the cost Hamiltonian and one application of the mixer Hamiltonian. We restrict to $p=1$, corresponding to the shallowest nontrivial QAOA circuit with two variational parameters, in order to remain within the NISQ-relevant shallow-depth regime.
@@ IV.3 Limitations
Our experiments were restricted to graphs of order four and five due to both hardware and algorithmic constraints. In the QUBO formulation, the number of logical qubits scales as $O(n^{2})$ with the number of vertices $n$, making even moderately larger instances infeasible on current NISQ devices with limited qubit counts, connectivity, and coherence times. Classical simulation is likewise hindered by the exponential growth of the Hilbert space. Further limitations arose from algorithmic overhead. The VQE, even with efficient optimizers such as SPSA, is challenging to scale to larger graphs, and remains resource-intensive. For these reasons, we confined our study to small graphs in order to establish proof-of-principle results. This choice enabled us to demonstrate energy clustering for isomorphic graph families under controlled conditions, while leaving the extension to larger and more complex graphs to future work.
However, we emphasize that the presence of energy clustering does not imply the existence of a strict energy threshold capable of separating isomorphic from non-isomorphic graph pairs. In practice, we observe substantial overlap in the low-energy spectra of isomorphic and non-isomorphic instances, particularly at shallow circuit depth. This overlap is a consequence of the highly degenerate structure of the isomorphism Hamiltonian and the fact that many low-energy states correspond to infeasible assignments that violate permutation constraints. Accordingly, when two data points in Fig. 1 share the same or nearly the same energy, this indicates convergence to energetically similar configurations, not a positive identification of graph isomorphism. To assess how fundamentally limited energy-based discrimination is within our variational framework, we further employed a suite of classical machine learning models trained on quantum-derived features, including final variational energy, optimization depth, and energy descent characteristics. These models—ranging from linear classifiers to expressive ensemble methods—represent a strong post-processing baseline for extracting structure from the variational outputs.
Despite this, classification performance remains modest, with no model achieving reliable separation between isomorphic and non-isomorphic graph pairs. Importantly, the failure of these models cannot be attributed to insufficient classifier expressivity, but rather indicates that the variational signal itself lacks sufficient discriminatory information. This suggests that the observed overlap in low-energy spectra reflects an intrinsic limitation of shallow variational encodings of the isomorphism Hamiltonian, rather than a shortcoming of any particular post-processing strategy. Progress in compact QUBO encodings, problem-tailored ansätze, and error mitigation techniques will be necessary to overcome these current limitations.
V Statistical Post-Processing Framework
The output of a Variational Quantum Algorithm (VQA) for the Graph Isomorphism problem is a stochastic energy expectation value, $\langle H_{C}\rangle$, rather than a binary decision variable. In an ideal, noise-free environment, isomorphic pairs would converge to a ground state energy of exactly zero, while non-isomorphic pairs would yield a strictly positive residual. However, in the NISQ era, this separation is obscured by stochastic noise, spectral crowding, and imperfect optimization convergence.
To rigorously distinguish the quantum signal from these noise sources, raw energy values are insufficient. We therefore employ a multi-tiered analysis suite—ranging from dynamic thresholding to manifold learning—to map the decision boundaries of the quantum solver and quantify the distinguishability of the distributions.
VI Results and discussion
VI.1 QAOA Energy Landscapes
The energies calculated correspond to the minimum eigenvalues of the problem Hamiltonian derived via the MinimumEigenOptimizer. The distributions of these energies for graph orders $N=4$ and $N=5$ are presented in Fig.˜1.
Consistent with Ising-based formulations of graph isomorphism Bonati (2014), we observe that isomorphic graph pairs tend to concentrate around lower variational energy values for small system sizes. For $N=4$, this separation is visually apparent. However, as the number of vertices increases, the ability to discriminate graph classes based solely on the lowest observed energies rapidly degrades. In particular, for $N=5$, we find that the lowest-energy samples obtained by shallow QAOA frequently correspond to infeasible assignments that violate the permutation (bijection) constraints, despite achieving near-optimal objective values.
This phenomenon is a direct consequence of the geometry of the feasible set and the structure of penalty-based encodings. The set of valid permutation matrices constitutes only $N!$ configurations within a Hilbert space of dimension $2^{N^{2}}$, yielding a vanishingly small feasible fraction
[…]
so that low-energy states are overwhelmingly dominated by infeasible superpositions unless the variational ansatz explicitly enforces permutation structure. Penalty Hamiltonians introduce large degenerate manifolds of near-feasible configurations, leading shallow-depth variational optimization to favor smooth constraint violations over discrete feasibility.
While Szegedy Szegedy (2019) showed that QAOA expectation values are invariant under graph isomorphisms at fixed depth, this invariance does not imply completeness for graph discrimination, nor does it guarantee that the lowest-energy states encode valid isomorphisms. At shallow depth, the locality horizon of QAOA limits its sensitivity to global bijection constraints, causing distinct non-isomorphic graphs to exhibit overlapping low-energy spectra once feasibility is taken into account. Consequently, discrimination based on minimum energy alone becomes unreliable beyond the smallest instances, even when apparent clustering persists at higher energy scales.
Importantly, the final energy value represents only the endpoint of a variational optimization trajectory. Two graph instances may converge to similar final energies while exhibiting qualitatively different optimization dynamics, such as convergence speed, oscillatory behavior, or energy drop magnitude across iterations. This observation suggests that additional information encoded in the optimization process itself may carry discriminative structure that is not captured by the final energy alone.
To operationalize this insight, we construct a low-dimensional quantum-derived feature vector
$\mathbf{x}=\bigl[N,\;E_{\mathrm{final}},\;\text{steps},\;\Delta E\bigr],$ (8)
where $N$ is the number of vertices, $E_{\mathrm{final}}$ is the terminal variational energy, steps denotes the number of optimizer iterations, and $\Delta E$ is the total energy decrease over the optimization trajectory. We then benchmark a suite of standard classical classifiers on this feature set to assess whether trajectory-level information improves discrimination between isomorphic and non-isomorphic graph pairs.
Specifically, we evaluate seven models: Logistic Regression, SVM with RBF kernel, Random Forest, Gradient Boosting, XGBoost, LightGBM, and a Multi-Layer Perceptron (MLP). Figure 2 summarizes their classification performance, while the corresponding confusion matrices are shown in Figure 3 to highlight the dominant error modes.
Figure 2: Model Performance Comparison. Tree-based ensemble methods (e.g., Random Forest and Gradient Boosting) consistently outperform linear classifiers, indicating that the decision boundary induced by the quantum-derived feature space is inherently non-linear.
Figure 3: Confusion Matrix Grid. Confusion matrices for each classifier, illustrating the relative prevalence of Type I (false positive) and Type II (false negative) errors.
Quantitative performance metrics for all models are reported in Table 1. Consistent with the trends observed in Figure 2, ensemble-based methods achieve the highest accuracy and AUC, with Random Forest performing best overall. In contrast, linear models perform near chance level, reinforcing the conclusion that discrimination relies on non-linear interactions among trajectory-level features rather than simple thresholding on final energy.
Table 1: Model Performance Ranking
Model Accuracy AUC
Random Forest 0.7500 0.7917
Gradient Boosting 0.7381 0.7905
XGBoost 0.6429 0.7196
Neural Network 0.6071 0.6424
SVM (RBF) 0.5952 0.6094
LightGBM 0.5952 0.6745
Logistic Regression 0.5119 0.5966
@@ VI.1 QAOA Energy Landscapes
Figure 4: Energy response under controlled graph reduction. (a) A sequence of graphs obtained by systematic edge removal from the five-node complete graph. For each reduced graph, a copy of the same graph is used, yielding strictly isomorphic pairs. (b) Final variational energies obtained from QAOA for each isomorphic pair. As edges are removed, the Hamiltonian energy shifts accordingly, while remaining tightly clustered near the minimum for each isomorphism class. This experiment isolates the dependence of the energy landscape on graph structure rather than graph labeling.
VI.2 VQE as a variational probe of the isomorphism Hamiltonian
The QAOA results discussed above characterize the behavior of a shallow-depth variational circuit applied to the graph isomorphism Hamiltonian. To further probe the structure of the same Hamiltonian, we additionally study the Variational Quantum Eigensolver (VQE), which employs a deeper and more expressive ansatz together with a generic variational optimization strategy.
For the small instances considered here, VQE reliably converges to the ground-state energy of the encoded Hamiltonian. In particular, for isomorphic graph pairs, the optimized objective value approaches zero, while non-isomorphic pairs converge to strictly positive values, consistent with the construction of the QUBO formulation. Importantly, this agreement should not be interpreted as evidence of efficient graph discrimination: rather, it confirms that the Hamiltonian correctly encodes isomorphism at the level of its ground state when the variational search is sufficiently expressive.
A notable distinction between QAOA and VQE lies in the scale and structure of the energies encountered. Due to its deeper ansatz and larger parameter space, VQE explores a much broader region of Hilbert space, leading to energy values that are typically orders of magnitude larger than those observed in shallow QAOA runs. This difference reflects improved approximation of the true ground state, rather than a fundamentally different encoding of graph structure.
Despite these differences, both algorithms exhibit a common qualitative trend: energies associated with isomorphic graph families cluster tightly, while non-isomorphic instances populate higher-energy regions of the spectrum. This clustering behavior therefore arises from the structure of the isomorphism Hamiltonian itself, rather than from a specific choice of ansatz or optimizer.
The increased expressivity of VQE comes at a substantial computational cost. In our implementation, VQE employed a TwoLocal ansatz optimized using the SPSA algorithm over up to 600 iterations, requiring repeated circuit evaluations at each step. Even for four-node graphs, this resulted in significantly longer runtimes and higher resource consumption compared to QAOA. Figure 5 illustrates representative convergence traces for multiple VQE runs on four-node graph pairs, showing that energy values stabilize only after several hundred optimization steps.